Show that (~π’‘βˆ¨π’’)∧(π’‘βˆ§~𝒒) is a contradiction



1
Expert's answer
2022-06-01T17:53:27-0400

"\left( {\~p \vee q} \right) \wedge \left( {p \wedge \~q} \right) = \left( {\~p \wedge p \wedge \~q} \right) \vee \left( {q \wedge p \wedge \~q} \right) = \left( {\left( {\~p \wedge p} \right) \wedge \~q} \right) \vee \left( {\left( {q \wedge \~q} \right) \wedge p} \right) = \left( {0 \wedge \~q} \right) \vee \left( {0 \wedge p} \right) = 0 \vee 0 = 0"

Q.E.D.